The ratio of lengths of two wires made up of the same material is 3:1 and the ratio of their radii is 1:3. The ratio of increament of lengths on account of suspending same weight will be ?

Grade:11

1 Answers

Sanju
6 years ago
We have Stress = Force/Area
Area, a = πr   [means area is proportional to r2]
Therefore, a1/a2 = (r1/r2)= (1/3)= 1/9
S1/S2 = [F/a1] /  [F/a2] = a2/a= 9/1    [Same weight means force is same and gets cancelled each other]
The ratio of increament of lengths on account of suspending same weight will be: 9:1